Please explain clearly and show all steps. Thank you. Solve the following first order equation 2x dy/dx = x+3y
Answer #1

atiy 22 21 dy In Let pa seda 20) e Consider the DE andy = x+3y dn 7 dy au dy ² + ( x w y - (2) Y = du -3 a = -zsta - Elux Int